OpenStudy (anonymous):

The following function defines a recursive sequence. f(0) = -3 f(1) = 6 f(n) = -2•f(n -1) - f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -3, 6, -9, -12, … -3, 20, -95, 480, … -3, 6, -9, 12, … -3, -20, -95, -480, …

OpenStudy (anonymous):

You know it's the first or third option from what you are given. \[f(2)=-2 \times f(1) - f(0) = (-2)(6)-(-3)=-9.\]. Can you find the last term in the sequence?

OpenStudy (anonymous):

-12?

OpenStudy (anonymous):

For the first term, n=0. So you need f(0), which is given to you. The second term, n=1, so f(1) is needed. This was also given to you. Then we calculated the third term, f(2). To calculate the last term, n=3, so you want to do \[f(3)=-2 \times f(2) - f(1)\] Is this -12?

OpenStudy (anonymous):

Negative times a negative equals a positive right?

OpenStudy (anonymous):

Yes. So the answer is?

OpenStudy (anonymous):

-2 x f(2) = 12

OpenStudy (anonymous):

oops soory

OpenStudy (anonymous):

got mixed with something else lol

OpenStudy (anonymous):

-2 x f(2) = 18 - f(1) = 12

OpenStudy (anonymous):

Correct :)

OpenStudy (anonymous):

can you help me with a couple other questions?

OpenStudy (anonymous):

Sure. What else do you have?

OpenStudy (anonymous):

For the function y = f(x), what is the ordered pair for the point on the graph when x = b - 2? (b - 2, f(b -2)) (x, f(b)) (x, b - 2) (b - 2, f(b)) I think its A

OpenStudy (anonymous):

Actually C i think would be correct

OpenStudy (anonymous):

Your first answer was correct. Can you justify why?

OpenStudy (anonymous):

That's the part im kinda confused on. I know that if x is replaced with b-2 it'll look like this f(b-2) but why is the y also b-2? Isn't the x only b-2?

OpenStudy (anonymous):

Each little point on the xy-plane is a coordinate of the form (x,y). Now the question tells you y = f(x) for the particular graph of f(x). So each coordinate on this graph is defined by ( x , f(x) )

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

So what is the coordinate when x = b-2?

OpenStudy (anonymous):

(b-2, f(b-2)

OpenStudy (anonymous):

Yes. For x = b-2, the point on the graph is ( b-2, f(b-2) )

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem!